OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) http://www.mathsisfun.com/algebra/circle-equations.html so equation of a circle is (x-a)^2 + (y-b)^2 = r^2 when r = radius and given the center point (a, b) so plug in a=5, b=6, r=3 into the equation :)

OpenStudy (anonymous):

Hello Thank you so much! Have a good day! @sleepyhead314